# Aligning Perception, Reasoning, Modeling and Interaction: A Survey on Physical AI

### v3

@@ E World Modeling
The previous sections have traced two key paths, namely Physical Perception and Physics Reasoning, towards AI systems that perceive and understand the law of physics. Physical perception embodies the end-to-end learning approach, extracting physical patterns and regularities directly from sensory observations., while physics reasoning encodes scientific knowledge into symbolic representations and theoretical frameworks Therefore, leveraging physical principles as constraints to construct and predict future states of the world has emerged as a natural idea, which has further evolved into a prominent research direction in frontier AI research, namely the rise of world models.
World Modeling represents the capability of AI systems to generate, understand and predict the dynamics of simulated environments, enabling a wide range of applications from video generation and scene reconstruction to autonomous planning and decision-making.
Models with world modeling ability have three fundamental advantages for current AI development: (1) reducing the need for supervised data, (2) closing the simulation-to-reality gap, and (3) enabling low-cost, interpretable prediction of future states. These advantages arise because world models enable agents to learn through interaction with simulated environments rather than relying on massive amounts of data, and they can provide more fine-grained sensory data for embodied agents [156]. In this section, we categorize litertures of world models into three dimensions: (1) Generative Models: They rely on gradient descent algorithm to learn underlying dynamics of the world from large scale database; (2) Physics-enhanced Approaches: They are constructed based on the principles and laws of physics in symbolic representation, combining the principle of physics that govern our physical world with neural networks to simulate changes in world states; (3) Benchmarks: They provide the metrics used to evaluate model’s capabilities in qualitative and quantitative estimation for world modeling.
@@ E.1 Generative Models
Image Generation. Image generation models transform abstract information into static visual representations. This process encompasses not only traditional lighting models like Phong Shading [157] and material rendering like Cook-Torrance [158], but more crucially utilizes Physically Based Rendering (PBR) through physical simulators like NVIDIA Isaac Sim [159] and Unity ML-Agents [160] to simulate the interactions between light rays and objects, thereby producing images with high fidelity in both visual and physical perspectives. Recent work in computer vision has demonstrated the effectiveness of PBR towards closing the sim-to-real gap, with applications ranging from depth sensor simulation  [161] to 6D object pose estimation  [162], where PBR-generated training data significantly outperformed traditional rendering approaches.
Under the world model framework, image synthesis establish the connection between internal physical state representations and external visual observations. Contemporary neural rendering approaches, especially those based on Neural Radiance Fields (NeRF), such as Ref-NeRF [163] and ENVIDR [164], allow the rendering process to produce both visually convincing images and physically consistent results. This transformation process from physical simulation to image synthesis effectively translates the internal state representations of world models into an observable and interpretable form for embodied agents, establishing a critical visual foundation for temporal modeling and dynamic prediction.
@@
Figure 6: Video Generations by GPT4Motion (Figure courtesy of [165]).
TABLE II: Performance of Video Generation Models on representative physical and world modeling benchmarks.
Model PhysicsIQ [166]()($\uparrow$) PhyGen [167]()($\uparrow$) VideoPhy [168] $(\uparrow)$ WorldModel Bench[169]()($\uparrow$)
Sora [170] 0.10 0.44 0.28 6.11
Pika [171] 0.13 0.44 0.29
CogVideoX [172] 0.45 0.49 7.31
LaVie [173] 0.36 0.41
Kling [174] 0.49 8.82
Modern video generation models, such as diffusion-based approaches like GPT4Motion [165] and VLIPP [175], are progressively constrained by physics laws, making the generated video sequences not only visually coherent but also physically plausible. For example, as illustrated in Figure 6, GPT4Motion is constrained by physics laws through physical simulators like PyBullet [176] by using GPT-4 to generate Blender scripts that simulate realistic physics dynamics, ensuring both visual consistency and adherence to physical laws across video frames, while VLIPP employs chain-of-thought reasoning in vision-language models to predict physically plausible motion trajectories that guide diffusion models to generate temporally coherent and physically accurate video sequences. DINO-Foresight [177] supports multiple future understanding tasks and avoids reconstructing irrelevant details of pixel based methods.
This physics-constrained video generation capability is crucial for building reliable world models, as it serves as an evaluation metric of world models’ understanding of environmental dynamics. As shown in Table II, recent video generation models demonstrate varying performance across physical and world modeling benchmarks, indicating their capability to capture environmental dynamics. Other benchmark uses physical experiments to evaluate how well a model can grasp the laws of physics LABEL:Zhang2025MorpheusBP. The ability to generate physically plausible video sequences indicates that the underlying world model has developed robust representations of environmental dynamics, making such generation capability a valuable benchmark for assessing world model quality.
@@
Traditional 3D reconstruction methods are typically based on geometric methods like SFM [178] and MVS [179], so they demonstrate strong geometric perception capabilities and often neglect the significance of physical reasoning, resulting in reconstructed scenes that violate fundamental physical principles.
Under the world model framework, 3D scene reconstruction should not only recover static geometric structures, but also understand and reconstruct the physical properties and dynamic behaviors among different objects within the scene. This requires reconstruction algorithms to possess not only geometric perception capabilities, but also physical reasoning capabilities. Modern scene reconstruction methods, particularly neural implicit representation-based approaches like PhyRecon [180] and IDR [181], are making progress in this direction. Not only can they reconstruct precise geometric structures, but also learn and encode physical properties such as density, stiffness, and friction coefficients through physical simulators like MuJoCo [182] and Warp [183]. This capability enables reconstructed scenes to be not only visually realistic but also accurate in physical behavior, which is precisely the key requirement for building actionable world models.
@@
The core challenge facing existing world models is insufficient out-of-distribution physical reasoning capability. While large-scale systems like GAIA-1 [184] and DINO-world [185] excel at in-distribution tasks and can generate realistic future scenarios, when encountering counterfactual reasoning, novel object interactions, or unfamiliar physical constraints, these models rely on "case-based" rather than "rule-based" generalization, leading to dramatic performance degradation [186]. Although early physics-based simulation methods can maintain consistency with real-world physics through predefined physics engines [26] [27], their reliance on manually designed parameterization limits their generalization capability to unmodeled physical effects.
@@
To address these challenges, researchers have proposed the concept of ’Neural Physics’, a hybrid framework that preserves fundamental physical laws while using neural networks to learn components that are traditionally difficult to model [187]. In this way, world models can perform accurate physical reasoning through the combination of physical principles and neural networks. Furthermore, they can create systems capable of bridging symbolic knowledge and embodied experience, representing a hallmark of human-level environmental understanding.
@@
(1) Neuro-symbolic integration directly embeds known physical laws into the neural network architecture to ensure strict compliance with fundamental physical principles, with the neural network solely responsible for learning uncertain components that are challenging to model accurately. For example, PhysORD [25] models vehicles as controlled Lagrangian systems by directly embedding the Euler-Lagrange equations as hard constraints within the network, where the neural network only needs to estimate quantities that are difficult to compute accurately, such as potential energy gradients and external forces. This approach ensures strict enforcement of fundamental physical laws like energy conservation, thereby maintaining good generalization capability even under data-scarce conditions.
(2) Physics-Structured Neural ODEs retains the mathematical structure of physical equations in the neural ODE framework, where neural networks parameterize various components of the physical equations to enable continuous-time domain modeling. For example, MoSim [188]decomposes the rigid body dynamics equations into predictor and corrector parts. The predictor strictly follows the mathematical structure of rigid body dynamics, with inertia matrices, gravity terms, and control forces parameterized through specialized neural network modules. The corrector uses standard residual networks to address complex phenomena such as friction and collisions that are challenging to model explicitly. Time integration is performed via neural ODE solvers, enabling the system to maintain both the inherent structure of physical equations and the learning capacity of neural networks.
(3) Differentiable Physics Engines employs a decoupled architecture in which neural networks specialize in predicting physical properties and interaction parameters from sensor data, with these predictions subsequently input into differentiable classical physics solvers for final system state computation. FusionForce [189] utilizes deep neural networks to predict terrain physical properties such as geometric shape, friction coefficients, and stiffness, along with robot-terrain contact forces from camera images and LiDAR data. These predictions are subsequently fed into a differentiable rigid body dynamics solver for robot motion trajectory computation. The end-to-end differentiable framework allows the model to refine terrain property predictions by backpropagating trajectory errors.
@@ E.3 Benchmarking Modeling Capability
(1) Fidelity evaluation focuses on the degree of similarity between model-generated content and real-world data. Image generation employs metrics such as FID [190] and CLIP Score [191], with evaluation conducted on datasets including COCO-30K [88] and ImageNet [89]. Video generation usually uses feature methods like I3D [192] , with evaluation conducted on Kinetics-400 [193].
@@
(2) Consistency evaluation examines the logical coherence of generated content under different conditions. This includes identity preservation tests on CelebA-HQ [194], action consistency tests on UCF-101 [195], and object state change evaluation on the Something-Something dataset [196].
@@
(3) Spatio-temporal coordination evaluation targets the spatial-temporal coordination capability of video generation. Commonly used benchmarks include Moving MNIST [197] motion prediction, BAIR Robot Pushing [198] object interaction evaluation, and Kinetics-700 action-scene coordination tests.
@@
(4) Physical plausibility assessment evaluates whether generated content adheres to physical laws, utilizing NYU Depth V2 [199] for image depth consistency tests, CLEVR-Video [200] for video physical reasoning tests, and the Real World Physics dataset for assessing gravity and collision prediction accuracy.
@@
Dynamic Modeling Evaluation. Evaluating the dynamics modeling capability of world models is crucial for assessing their understanding of physical laws and temporal evolution. Current evaluation frameworks can be classified by multiple dimensions: (1) general trajectory prediction: For moving objects, models must forecast future positions based on initial conditions and physical constraints [201, 202, 203, 204]; (2) deformable material simulation fidelity: Require modeling of elasticity and collision dynamics, particularly for complex materials like cloth and rubber [205, 206, 207, 208, 209]; and (3) fluid dynamics modeling: Test model ability to capture turbulence, viscosity, and flow patterns [210, 211, 212]. Some comprehensive benchmarks also provides standardized evaluation scenarios covering rigid-body collisions, soft-body deformations, and projectile motion, and is widely used to measure models’ short-term prediction accuracy and long-term stability [213, 214, 215, 216]. Critical metrics in these benchmarks include physical consistency (adherence to conservation laws), state transition accuracy (alignment between predicted and ground-truth trajectories), and generalization to novel physical configurations.
@@
Virtual Interaction Capability. The ability to accurately model virtual interactions is fundamental to evaluating world models’ understanding of cause-and-effect relationships and their capacity for counterfactual reasoning. Virtual environments like game engines provide an ideal crucible for testing the predictive accuracy and planning capabilities of an agent’s internal world model.The AIBIRDS competition exemplifies this approach, requiring agents to solve unseen levels in real-time through accurate simulation of complex physical interactions [217], thus testing both predictive accuracy and adaptive planning capabilities within physically embedded systems [218]. Some standard benchmarks such as Hi-Phy, PHYRE and PhyQ assess generalization across diverse interaction templates [219, 220, 221], while adversarial evaluations probe deeper understanding through deceptive scenarios that punish superficial heuristics and demand sophisticated multi-step reasoning [222]. Other works also demand that explicit physics simulators deliver interpretable, physics-aware decision support in complex interaction scenarios [223]. Ultimately, these virtual evaluating environments demonstrate that achieving robust physics intelligence requires more than just prediction, but involving human-like decision-making strategies such as knowing when to stop loss and restart a failed attempt [224], where this kind of trial-and-error learning fundamental to solving real-world challenges like marble runs [225].
@@ F Embodied Interaction
World models provide the possibility for modeling virtual physical scenarios, yet a substantial gap persists between simulation environments and the real world. Embodied intelligence, as an innovative technology that demands multi-dimensional reasoning and physical interaction capabilities, transcends the limitations of approaches that primarily focus on predictions within controlled spaces, progressively emerging as the ultimate paradigm for the development of future physics based AI models.
@@ F.1 Robotics
While conventional world models excel at predicting outcomes in simulated environments, robotics applications demand the critical transition from passive observation to active manipulation. This evolution is exemplified by the emergence of Vision-Language-Action (VLA) models [226] that bridge the gap between building internal world representations and changing external physical reality. Pioneering systems like Gato [227] and RT-1 [228] establishing the foundation for this paradigm shift.
Recent advances including $\pi_{0}$ [229], OpenVLA [230], and Gemini Robotics [231] further demonstrate how physics-enhanced AI with embodiment extends beyond simulation-based world models by requiring real-time adaptation to physical constraints, sensor noise, and the irreversible consequences of actions in the real world.
Figure 7: Overview of OpenVLA architecture (Figure used courtesy of [230]).
The transition from simulated world models to reality systems also introduces new challenges. First, continuous action generation highlights the inadequacy of discretized representations when confronting the smooth, continuous nature of reality, because robotics demands continuous control that respects physical constraints [232]. Models like $\pi_{0}$ address this through flow matching [233], generating physically plausible trajectories that maintain stability during contact-rich manipulation. This direction is further explored through diffusion policies [234] . Furthermore, cross platform generalization tests whether learned physical principles can transcend specific hardware implementations. As shown in Figure 7, OpenVLA’s approach of learning normalized action deltas from diverse datasets [235] demonstrates how embodied model must abstract beyond simulation specific assumptions to achieve morphology independent reasoning that generalizes across varied physical platforms. Third, real-world perceptual grounding requires models to construct accurate world representations from noisy, partial observations. Some works addresses this through enhanced multi-view correspondence understanding [231], object localization [236, 237] and physical properties recognition [238], capabilities essential for reasoning about physical interactions under perceptual uncertainty.

### v4

@@ E World Modeling
The previous sections have traced two key paths, namely Physical Perception and Physics Reasoning, towards AI systems that perceive and understand the law of physics. Physical perception embodies the end-to-end learning approach, extracting physical patterns and regularities directly from sensory observations, while physics reasoning encodes scientific knowledge into symbolic representations and theoretical frameworks.
After internalizing such physical knowledge, can models strengthen their understanding of governing principles of the world and thereby predict the future? World models, as an emerging line of research, offer a natural next step for physical AI systems by enabling a progression from understanding to modeling. World Modeling represents the capability of AI systems to generate, understand and predict the dynamics of simulated environments, enabling a wide range of applications from video generation and scene reconstruction to autonomous planning and decision-making.
Models with world modeling ability have three fundamental advantages for current AI development: (1) reducing the need for supervised data, (2) closing the simulation-to-reality gap, and (3) enabling low-cost, interpretable prediction of future states. These advantages arise because world models enable agents to learn through interaction with simulated environments rather than relying on massive amounts of data, and they can provide more fine-grained sensory data for embodied agents [understanding_or_pridiction].
Unlike physics reasoning LLMs, world models exhibit substantial shifts in their architectural design, training paradigms, and evaluation methodologies. In this section, we categorize litertures of world models into three dimensions: (1) Generative Models: They rely on gradient descent algorithm to learn underlying dynamics of the world from large scale database; (2) Physics-enhanced Approaches: They are constructed based on the principles and laws of physics in symbolic representation, combining the principle of physics that govern our physical world with neural networks to simulate changes in world states; (3) Benchmarks: They provide the metrics used to evaluate model’s capabilities in qualitative and quantitative estimation for world modeling.
@@ E.1 Generative Models
Image Generation. Image generation models transform abstract information into static visual representations. This process encompasses not only traditional lighting models like Phong Shading [Phong] and material rendering like Cook-Torrance [Cook-Torrance], but more crucially utilizes Physically Based Rendering (PBR) through physical simulators like NVIDIA Isaac Sim [isaacsim] and Unity ML-Agents [Unity-ML-Agents] to simulate the interactions between light rays and objects, thereby producing images with high fidelity in both visual and physical perspectives. Recent work in computer vision has demonstrated the effectiveness of PBR towards closing the sim-to-real gap, with applications ranging from depth sensor simulation  [stereovision-depth-sensors] to 6D object pose estimation  [BOP], where PBR-generated training data significantly outperformed traditional rendering approaches.
Under the world model framework, image synthesis establish the connection between internal physical state representations and external visual observations. Contemporary neural rendering approaches, especially those based on Neural Radiance Fields (NeRF), such as Ref-NeRF [Ref-NeRF] and ENVIDR [ENVIDR], allow the rendering process to produce both visually convincing images and physically consistent results. This transformation process from physical simulation to image synthesis effectively translates the internal state representations of world models into an observable and interpretable form for embodied agents, establishing a critical visual foundation for temporal modeling and dynamic prediction.
@@
Figure 6: Video Generations by GPT4Motion (Figure courtesy of [GPT4Motion]).
TABLE III: Performance of Video Generation Models on representative physical and world modeling benchmarks.
Model PhysicsIQ [Physics-IQ]()($\uparrow$) PhyGen [phygen]()($\uparrow$) VideoPhy [2024videophy] $(\uparrow)$ WorldModel Bench[worldmodelbench]()($\uparrow$)
Sora [opensora] 0.10 0.44 0.28 6.11
Pika [Pika] 0.13 0.44 0.29
CogVideoX [CogVideoX] 0.45 0.49 7.31
LaVie [wang2023lavie] 0.36 0.41
Kling [Kling] 0.49 8.82
Modern video generation models, such as diffusion-based approaches like GPT4Motion [GPT4Motion] and VLIPP [VLIPP], are progressively constrained by physics laws, making the generated video sequences not only visually coherent but also physically plausible. For example, as illustrated in Figure 6, GPT4Motion is constrained by physics laws through physical simulators like PyBullet [PyBullet] by using GPT-4 to generate Blender scripts that simulate realistic physics dynamics, ensuring both visual consistency and adherence to physical laws across video frames, while VLIPP employs chain-of-thought reasoning in vision-language models to predict physically plausible motion trajectories that guide diffusion models to generate temporally coherent and physically accurate video sequences. DINO-Foresight [Karypidis2024DINOForesightLI] supports multiple future understanding tasks and avoids reconstructing irrelevant details of pixel based methods.
This physics-constrained video generation capability is crucial for building reliable world models, as it serves as an evaluation metric of world models’ understanding of environmental dynamics. As shown in Table III, recent video generation models demonstrate varying performance across physical and world modeling benchmarks, indicating their capability to capture environmental dynamics. The ability to generate physically plausible video sequences indicates that the underlying world model has developed robust representations of environmental dynamics, making such generation capability a valuable benchmark for assessing world model quality.
@@
Traditional 3D reconstruction methods are typically based on geometric methods like SFM [SFM] and MVS [MVS], so they demonstrate strong geometric perception capabilities and often neglect the significance of physical reasoning, resulting in reconstructed scenes that violate fundamental physical principles.
Under the world model framework, 3D scene reconstruction should not only recover static geometric structures, but also understand and reconstruct the physical properties and dynamic behaviors among different objects within the scene. This requires reconstruction algorithms to possess not only geometric perception capabilities, but also physical reasoning capabilities. Modern scene reconstruction methods, particularly neural implicit representation-based approaches like PhyRecon [PhyRecon] and IDR [IDR], are making progress in this direction. Not only can they reconstruct precise geometric structures, but also learn and encode physical properties such as density, stiffness, and friction coefficients through physical simulators like MuJoCo [MuJoCo] and Warp [warp2022]. This capability enables reconstructed scenes to be not only visually realistic but also accurate in physical behavior, which is precisely the key requirement for building actionable world models.
@@
The core challenge facing existing world models is insufficient out-of-distribution physical reasoning capability. While large-scale systems like GAIA-1 [GAIA-1] and DINO-world [DINO-world2025] excel at in-distribution tasks and can generate realistic future scenarios, when encountering counterfactual reasoning, novel object interactions, or unfamiliar physical constraints, these models rely on "case-based" rather than "rule-based" generalization, leading to dramatic performance degradation [physical-law-perspective2024]. Although early physics-based simulation methods can maintain consistency with real-world physics through predefined physics engines [differentiable-stokes-flow2020] [Scalable-differentiable-physic2020], their reliance on manually designed parameterization limits their generalization capability to unmodeled physical effects.
@@
To address these challenges, researchers have proposed the concept of ’Neural Physics’, a hybrid framework that preserves fundamental physical laws while using neural networks to learn components that are traditionally difficult to model [neural_physics_2025]. In this way, world models can perform accurate physical reasoning through the combination of physical principles and neural networks. Furthermore, they can create systems capable of bridging symbolic knowledge and embodied experience, representing a hallmark of human-level environmental understanding.
@@
(1) Neuro-symbolic integration directly embeds known physical laws into the neural network architecture to ensure strict compliance with fundamental physical principles, with the neural network solely responsible for learning uncertain components that are challenging to model accurately. For example, PhysORD [PhysORD2024] models vehicles as controlled Lagrangian systems by directly embedding the Euler-Lagrange equations as hard constraints within the network, where the neural network only needs to estimate quantities that are difficult to compute accurately, such as potential energy gradients and external forces. This approach ensures strict enforcement of fundamental physical laws like energy conservation, thereby maintaining good generalization capability even under data-scarce conditions.
(2) Physics-Structured Neural ODEs retains the mathematical structure of physical equations in the neural ODE framework, where neural networks parameterize various components of the physical equations to enable continuous-time domain modeling. For example, MoSim [MoSim2025]decomposes the rigid body dynamics equations into predictor and corrector parts. The predictor strictly follows the mathematical structure of rigid body dynamics, with inertia matrices, gravity terms, and control forces parameterized through specialized neural network modules. The corrector uses standard residual networks to address complex phenomena such as friction and collisions that are challenging to model explicitly. Time integration is performed via neural ODE solvers, enabling the system to maintain both the inherent structure of physical equations and the learning capacity of neural networks.
(3) Differentiable Physics Engines employs a decoupled architecture in which neural networks specialize in predicting physical properties and interaction parameters from sensor data, with these predictions subsequently input into differentiable classical physics solvers for final system state computation. FusionForce [fusionforce2025] utilizes deep neural networks to predict terrain physical properties such as geometric shape, friction coefficients, and stiffness, along with robot-terrain contact forces from camera images and LiDAR data. These predictions are subsequently fed into a differentiable rigid body dynamics solver for robot motion trajectory computation. The end-to-end differentiable framework allows the model to refine terrain property predictions by backpropagating trajectory errors.
@@ E.3 Benchmarking Modeling Capability
(1) Fidelity evaluation focuses on the degree of similarity between model-generated content and real-world data. Image generation employs metrics such as FID [FID] and CLIP Score [Clipscore], with evaluation conducted on datasets including COCO-30K [coco] and ImageNet [Imagenet]. Video generation usually uses feature methods like I3D [I3D] , with evaluation conducted on Kinetics-400 [Kinetics-400].
@@
(2) Consistency evaluation examines the logical coherence of generated content under different conditions. This includes identity preservation tests on CelebA-HQ [CelebA-HQ], action consistency tests on UCF-101 [Ucf101], and object state change evaluation on the Something-Something dataset [Something-Something].
@@
(3) Spatio-temporal coordination evaluation targets the spatial-temporal coordination capability of video generation. Commonly used benchmarks include Moving MNIST [MovingMNIST] motion prediction, BAIR Robot Pushing [BAIR] object interaction evaluation, and Kinetics-700 action-scene coordination tests.
@@
(4) Physical plausibility assessment evaluates whether generated content adheres to physical laws, utilizing NYU Depth V2 [NYU-Depth-V2] for image depth consistency tests, CLEVR-Video [CLEVR] for video physical reasoning tests, and the Real World Physics dataset for assessing gravity and collision prediction accuracy.
@@
Dynamic Modeling Evaluation. Evaluating the dynamics modeling capability of world models is crucial for assessing their understanding of physical laws and temporal evolution. Current evaluation frameworks can be classified by multiple dimensions: (1) general trajectory prediction: For moving objects, models must forecast future positions based on initial conditions and physical constraints [Bullinger20173DTRA, PDDL, Zhang2024M3BenchBW, Zhu2023BenchmarkingSV]; (2) deformable material simulation fidelity: Require modeling of elasticity and collision dynamics, particularly for complex materials like cloth and rubber [Wang2018NeuralML, Liang2019DifferentiableCSA, Mianroodi2021LosslessMCA]; and (3) fluid dynamics modeling: Test model ability to capture turbulence, viscosity, and flow patterns [predictivemodels, Zhao_2024, liquidsgraphnetworks]. Some comprehensive benchmarks also provides standardized evaluation scenarios covering rigid-body collisions, soft-body deformations, and projectile motion, and is widely used to measure models’ short-term prediction accuracy and long-term stability [Physion, PhysionPlus, ContPhy, PhyWorldBench]. Critical metrics in these benchmarks include physical consistency (adherence to conservation laws), state transition accuracy (alignment between predicted and ground-truth trajectories), and generalization to novel physical configurations.
@@
Virtual Interaction Capability. The ability to accurately model virtual interactions is fundamental to evaluating world models’ understanding of cause-and-effect relationships and their capacity for counterfactual reasoning. Virtual environments like game engines provide an ideal crucible for testing the predictive accuracy and planning capabilities of an agent’s internal world model.The AIBIRDS competition exemplifies this approach, requiring agents to solve unseen levels in real-time through accurate simulation of complex physical interactions [AIBIRDS], thus testing both predictive accuracy and adaptive planning capabilities within physically embedded systems [physicallyembeddedplanningproblems]. Some standard benchmarks such as Hi-Phy, PHYRE and PhyQ assess generalization across diverse interaction templates [Hi-Phy, phyre, Phy-Q], while adversarial evaluations probe deeper understanding through deceptive scenarios that punish superficial heuristics and demand sophisticated multi-step reasoning [deceptive]. Other works also demand that explicit physics simulators deliver interpretable, physics-aware decision support in complex interaction scenarios [CueTip]. Ultimately, these virtual evaluating environments demonstrate that achieving robust physics intelligence requires more than just prediction, but involving human-like decision-making strategies such as knowing when to stop loss and restart a failed attempt [usingrestartheuristicsimprove], where this kind of trial-and-error learning fundamental to solving real-world challenges like marble runs [learningexplorationstrategiessolve].
@@ F Embodied Interaction
World models provide the possibility for modeling virtual physical scenarios, yet a substantial gap persists between simulation environments and the real world. This gap highlights an essential limitation: modeling alone is insufficient for mastering the dynamics of real physical processes. Embodied intelligence addresses this limitation by requiring agents to perform multi-dimensional reasoning and interact with their physical surroundings. As a result, it moves beyond approaches that operate solely within controlled or virtual spaces and is increasingly viewed as a key paradigm for advancing future physics-based AI.
@@ F.1 Robotics
While conventional world models excel at predicting outcomes in simulated environments, robotics applications demand the critical transition from passive observation to active manipulation. This evolution is exemplified by the emergence of Vision-Language-Action (VLA) models [VLA] that bridge the gap between building internal world representations and changing external physical reality. Pioneering systems like Gato [ReedEtAl2022_Gato] and RT-1/2 [BrohanEtAl2022_RT1, Brohan2023RT2VM] establishing the foundation for this paradigm shift.
Recent advances including $\pi_{0}$ [BlackEtAl2024_pi0], OpenVLA [KimEtAl2024_OpenVLA], and Gemini Robotics [DeepMind2024_GeminiRobotics] further demonstrate how physics-enhanced AI with embodiment extends beyond simulation-based world models by requiring real-time adaptation to physical constraints, sensor noise, and the irreversible consequences of actions in the real world. Figure 7 illustrates primary paradigms for current VLA systems. Both fusion-based (late/early) and dual-system architectures reflect the importance of integrating fast physical control with deliberative reasoning.
[…]
Figure 7: Architecture patterns for vision-language-action models. The shift from Late Fusion single-stream architectures to Dual-System designs reflects the integration of physical reasoning into VLAs.
The transition from simulated world models to reality systems also introduces new challenges. First, continuous action generation highlights the inadequacy of discretized representations when confronting the smooth, continuous nature of reality, because robotics demands continuous control that respects physical constraints [Safety-Critical]. Models like $\pi_{0}$ address this through flow matching [lipman2023_flow_matching], generating physically plausible trajectories that maintain stability during contact-rich manipulation. This direction is further explored through diffusion policies [ChiEtAl2023_DiffusionPolicy] . Furthermore, cross platform generalization tests whether learned physical principles can transcend specific hardware implementations. OpenVLA’s approach of learning normalized action deltas from diverse datasets [BrohanEtAl2022_OpenX] demonstrates how embodied model must abstract beyond simulation specific assumptions to achieve morphology independent reasoning that generalizes across varied physical platforms. Third, real-world perceptual grounding requires models to construct accurate world representations from noisy, partial observations. Some works addresses this through enhanced multi-view correspondence understanding [DeepMind2024_GeminiRobotics], object localization [LingEtAl2023_ActiveLearning, SeoEtAl2023_HybridControl] and physical properties recognition [WuEtAl2023_ObjectProperty], capabilities essential for reasoning about physical interactions under perceptual uncertainty.